OpenStudy (chosenmatt):

What is the exact circumference of the circle? A.48 inches B.32 inches C.16 inches D.8 inches

OpenStudy (igreen):

Formula: \(V = \pi d\)

OpenStudy (igreen):

Where 'd' is the diameter

OpenStudy (igreen):

So just plug in 16 for 'd'.

pooja195 (pooja195):

\[C=2πr \] wouldnt this be the formula?

OpenStudy (chosenmatt):

so 32

OpenStudy (igreen):

Yes, but 2 times the radius is the same as the diameter. \(C = 2 \pi r\) or \(C = \pi d\)

OpenStudy (anonymous):

is the same

pooja195 (pooja195):

ah ok :) continue on then

OpenStudy (igreen):

How did you get 32? @chosenmatt

OpenStudy (chosenmatt):

16 times 2

OpenStudy (igreen):

Where did you get 2 from?

OpenStudy (igreen):

\(\color{blue}{\text{Originally Posted by}}\) @iGreen Formula: \(V = \pi d\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (chosenmatt):

xD idk lol i guess i was just trying something

OpenStudy (igreen):

x Just plug in 16 for 'd'..

OpenStudy (igreen):

*xD

OpenStudy (chosenmatt):

so C

OpenStudy (igreen):

Yeah, if the option is \(16 \pi\)..is it?

OpenStudy (chosenmatt):

yes

OpenStudy (igreen):

Yep, then that's our answer.

OpenStudy (chosenmatt):

couldnt put pie tho :/

OpenStudy (igreen):

'Exact', so I guessed they all had 'pi' next to it.

OpenStudy (chosenmatt):

i mean pi xD

OpenStudy (igreen):

``` \(\pi\)

OpenStudy (chosenmatt):

:} i want pie now xD lol

OpenStudy (igreen):

xD Don't worry, pi day is in 10 days :)